Find the inverse Laplace transforms of the following functions:

(a)

s + 3 /(s^2 + 6s + 13)

(b)

2s + 3/(s^2+6s+13)


1
Expert's answer
2021-07-19T16:58:20-0400

(a)


"L^{-1}(\dfrac{s+3}{s^2+6s+13})=L^{-1}(\dfrac{s-(-3)}{(s-(-3))^2+2^2})"

"=e^{-3t}\cos(2t)"


(b)


"L^{-1}(\dfrac{2s+3}{s^2+6s+13})=L^{-1}(\dfrac{2(s-(-3))-\dfrac{3}{2}(2)}{(s-(-3))^2+2^2})"

"=2L^{-1}(\dfrac{s-(-3)}{(s-(-3))^2+2^2})"

"-\dfrac{3}{2}L^{-1}(\dfrac{2}{(s-(-3))^2+2^2})"

"=2e^{-3t}\cos(2t)-\dfrac{3}{2}e^{-3t}\sin(2t)"